OpenStudy (anonymous):

How do you derive ohm's law from the current density and electric field within a wire of uniform thickness???? Any hints??

OpenStudy (anonymous):

If you have the electric field and the thickness of the wire, you should be able to dig out electric flux; use that and the current density to derive it. HINT: You've got current...can you find resistance? Note: I'm skipping a step or two in here intentionally, so you can figure the rest out.

OpenStudy (anonymous):

we know, \[J=\sigma E\]\[where \sigma=\mu n e\] now we knoe J=I/A, E=V/d I/A=sigma*V/d \[V=\frac{d}{A \sigma}I\] \[\frac1{\sigma}=\rho\] u get \[V=\rho\frac lA I\]that is our Ohm's law

OpenStudy (anonymous):

good

OpenStudy (nilankshi):

good ans